# Physics-Informed Detection of Friction Anomalies in Satellite Reaction Wheels

### v1

@@ III-B Friction Switching System
An FSS represents a particular phenomenon inside the bearing of the RWA, such as the distribution of lubricant, the relative position of the inner and outer races, or any other aspect that may affect the dry friction. In these systems, the variable $q$ represents the current configuration of the system, $\tau$ is how long the system has been in that configuration, and $f$ is the friction generated by the system. Note that the configuration only takes discrete values.
These systems are indexed with the variable $s$, and the evolution of a system $s$ is relatively simple. At every step, the system may jump to a contiguous configuration $q$ with $P_{s}^{\mathrm{t}}(q,\tau)$ or stay the same. If it does not jump, the configuration does not change, and $\tau$, the number of steps since the last jump, increases by one. If it jumps, the configuration $q$ and the friction $f$ produced by the FSS change, and $\tau$ goes back to 0. A visualization of a system FSS can be seen in Figure 3, and the dynamics are given formally by
$\left[\begin{matrix}q_{k+1}\\ f_{k+1}\\ \tau_{k+1}\end{matrix}\right]$ $=\begin{cases}\left[\begin{matrix}q_{k}&f_{k}&\tau_{k}+1\end{matrix}\right]^{\top}&\text{with prob. }P_{s}^{\mathrm{t}}(q_{k},\tau_{k}),\\ \left[\begin{matrix}q^{\prime}&f^{\prime}&0\end{matrix}\right]^{\top}&\text{otherwise}.\end{cases}$ (5)
Two probability functions affect the behavior of the FSS. The first is the transition probability $P_{s}^{t}(q,\tau)$, which depends both on the current configuration and on how long the system has spent in it. By including $q$, we can consider systems that spend different lengths of time in various configurations, and by including $\tau$, we are not limited to exponential residence times, as it is common in stochastic Markov chains [43].
The second probability function is $P_{s}^{\mathrm{qf}}(q^{\prime},f^{\prime})$, which determines the new configuration and friction ($q^{\prime}$ and $f^{\prime}$ in (5)) the system takes after it has transitioned. Since the friction depends only on the new configuration, this can be written as
$(q^{\prime},f^{\prime})\sim P_{s}^{\mathrm{qf}}(q^{\prime},f^{\prime}|q)=P_{s}^{\mathrm{f}}(f^{\prime}|q^{\prime})P_{s}^{\mathrm{q}}(q^{\prime}|q).$ (6)
We set some assumptions on $P_{s}^{\mathrm{f}}(f^{\prime}|q^{\prime})$ and $P_{s}^{\mathrm{q}}(q^{\prime}|q)$. For the latter, the FSS can only jump between contiguous states, that is, $P_{s}^{\mathrm{q}}(q^{\prime}|q)=0$ if $|q^{\prime}-q|>1$. For the friction: the probability functions associated with different $q$ can not overlap, and the range of values of $f^{\prime}$ must increase with increasing $q$. Figure 3 shows an example of valid friction probabilities on the right side.
To characterize a FSS, we need four elements: its number of configurations $q_{\text{max}}$, its transition probability function $P_{s}^{\mathrm{q}}(q^{\prime}|q)$, its jumping probability function $P_{s}^{\mathrm{t}}(q,\tau)$ for every configuration $q$ and the friction probability function $P_{s}^{\mathrm{f}}(f|q)$, also for every $q$. An example of the friction produced by an RWA modeled with two FSS is shown in Figure 4, where its similarity to the examples shown in Figure 2 can be noted.
@@
Figure 3: Evolution of the state and friction of a switching system. The left nodes show the different configurations an FSS can take (3 in this example) and the possible transitions between them. The right side shows the friction distributions of each configuration, together with a possible trajectory of the friction value generated by the FSS as it jumps between configurations.
Example 1 (Short term friction changes)
@@
$P_{1}^{\mathrm{f}}(f^{\prime}|q^{\prime})$ $=\begin{cases}10\quad&\text{if }0.3q^{\prime}<f<0.3q^{\prime}+0.1\\ 0\quad&\text{otherwise}\end{cases}$ (7a)
$P_{1}^{\mathrm{t}}(\tau|q)$ $=\begin{cases}1/(200-\tau)\quad&\text{if }q=0\\ 1/(2000-\tau)\quad&\text{if }q=1\end{cases}$ (7b)
Example 2 (Longer term friction changes)
We want to model long-term variations of the friction switching between three different configurations, with the same transition probabilities. This can be due to, for example, changes in the angle of the cage inside the bearing. The probability functions are given by:
$P_{2}^{\mathrm{f}}(f^{\prime}|q^{\prime})$ $=\begin{cases}1.5\quad&\text{if }0.5q^{\prime}<f^{\prime}<0.5q^{\prime}+0.4\\ 0\quad&\text{otherwise}\end{cases}$ (8a)
$P_{2}^{\mathrm{t}}(x,\tau)$ $=\begin{cases}\frac{1}{30000-\tau}\quad&\text{if }x>10000,\\ 0\quad&\text{otherwise}.\end{cases}$ (8b)
We show a simulation of the friction system proposed in Equations (4a) and (4b) with the switching system using the probability functions described above and the parameters $f^{\bar{\mathrm{d}}}=1$ and $f^{\mathrm{v}}=1$. The evolution of the friction components in time, along with the total dry friction coefficient, is shown in Figure 4.
Figure 4: Visualization of the two friction systems described in Examples 1 and 2. The first three plots show the evolution of the three state variables of the system. The fourth plot shows the resulting total dry friction, and the fifth the total friction after adding a viscous component.
@@
In the paper, we assume knowledge of $q_{\text{max}}$, $P_{s}^{\mathrm{t}}(x,\tau)$ and $P_{s}^{\mathrm{q}}(q^{\prime}|q)$ for every FSS in the RWA we are analyzing. Such knowledge may come from historical data, expert knowledge, or both. On the other hand, we do not know the friction probability function $P_{s}^{\mathrm{f}}(f^{\prime}|q^{\prime})$, which might be affected by an anomaly.
@@ III-C Anomaly Modeling
To conclude this section, we define how the possible anomalies in the RWA are represented in the proposed model. As mentioned in Section I, there are several possible anomalies, and all of them affect the resulting friction experienced in the bearing of the RWA.
We consider each anomaly to relate to one and only one component of friction. For instance, the dry friction anomaly $\theta^{\mathrm{d}}$ affects the base dry friction coefficient, while the viscous friction anomaly $\theta^{\mathrm{v}}$ affects the viscous friction coefficient.
For the case of the switching systems, each FSS present in the RWA can be affected by one anomaly, represented by $\theta_{s}$, where $s$ is the index of the affected FSS. In this model, we consider that the effect of the anomaly manifests itself only in the friction distribution function, that is, in the values of the friction the system produces for every configuration. This relationship can be expressed mathematically as:
@@
Although the probability functions are unknown, they can be estimated using a dataset $\mathcal{D}$, hence the need for a hybrid approach.
IV Algorithm Description
The proposed anomaly detection algorithm uses the model developed in Section III to analyze the window of measurements under investigation and correctly diagnose the state of the RWA. As explained in Section III, the possible anomalies of the RWA affect the values of friction that each component (base dry, viscous and the FSSs) produces. The objective, then, is to extract the relevant information from the data to do the classification.
The main challenge in the model presented in the previous section is the presence of several FSSs, each producing sudden changes in friction at random moments. It is then fundamental to keep track of these changes to detect and categorize the different anomalies.
For this purpose, the algorithm comprises four stages, each with a clear responsibility in the complete process. The first stage, changepoint detection, determines the time-steps in which an abrupt friction change has occurred. The second stage is the friction estimation, which determines the dry and viscous friction coefficients before and after every changepoint, that is, $f_{k}^{\mathrm{d}}$ and $f^{\mathrm{v}}$ in (4a). The different friction coefficients can then be used to calculate how much the friction changes at every changepoint. The third stage is the maximum likelihood assignment, which determines which FSS has produced each changepoint by maximizing the likelihood of the observed data. With this information we can distinguish the individual contributions $f^{\bar{\mathrm{d}}}$ and $f_{k}^{s}$ in (4b). Finally, once the distribution of friction values of dry friction, viscous friction, and every switching system is known, the anomaly classification stage produces an estimated status $\hat{\theta}$ using an SVM trained on the dataset provided. A diagram illustrating the stages of the algorithm, along with the data flow between them, is shown in Figure 5.
@@ IV Algorithm Description
Figure 5: Diagram of the complete algorithm presented. The notation is explained along Section V.
@@ IV-A Changepoint Detection
The process of obtaining the regions of change of friction in the signal is divided into two parts. First, we compute a metric that estimates the probability that the friction has changed for every step $k$ in the data. Then, we find local maxima of this metric that are above a certain threshold. These maxima are our estimation of where the changepoints are located.
@@
with $C$ being a constant that gets canceled in the computation of the GLR, and $W_{\mathrm{b}}$ a parameter used to penalize large deviations of the viscous friction from the values that we expect to find in the data.
@@
Like the Log-Likelihood Ratio (LLR), the metric computes the log-likelihood of two different hypotheses: a jump taking place at $k$, or no jump at all. One of the issues with the LLR is the need to consider the variables involved in both hypotheses, that is, the dry and viscous friction coefficients. This problem can be addressed through marginalization, leading to the Marginalized Likelihood Ratio or through maximization for the parameters, as we do in this problem, leading to the Generalized Likelihood Ratio [44].
@@
The GLR was developed for cases where there is only one changepoint to detect, but there are several in our problem. Rather than computing the likelihood for the whole data, this issue is addressed by computing it in a window of $w$ points around the hypothetical changepoints. The constant $w$ is a hyperparameter of the algorithm selected from historical data, on the condition that this window size is smaller than the minimum distance between changepoints expected in the data.
@@
The last element in the metric is the use of a prior for the viscous friction $P(f^{\mathrm{v}})$. The purpose of this element is to avoid obtaining parameters in the optimization that fit the data well but are not realistic. As we will see in Section V, a high penalization constant $W_{\textrm{b}}$ can improve the algorithm’s effectiveness when the window size $w$ is small. The base viscous friction coefficient $\bar{f}^{\mathrm{v}}$ is obtained from historical data and, as it will be explained also in Section V, does not need to be very accurate to fulfill its objective.
The purpose of this last term is to penalize parameters that deviate too much from the expected values. Its tuning is not trivial, so a parameteric study is conducted in the results section. The base viscous friction, $\tilde{f}^{\mathrm{v}}$, is defined using historical data, and should approximate the expected value of this parameter.
@@
Once the GLR has been computed, we must determine which points correspond to changepoints. The original GLR algorithm solves this problem by taking the point $k$ at which the metric is maximized. In our case, however, we must take into account that there are several changepoints in the complete dataset. Therefore, we first find windows of contiguous measurements where the GLR is above a threshold $\text{GLR}_{\text{thr}}$ and, for each window, we designate the maximum among those points to be a changepoint $k^{\text{cp}}$. The number of changepoints will equal the number of windows above the threshold found.
@@
The results obtained by this stage of the algorithm depend on how its main hyperparameter, the threshold, is set. Under the assumption that the dry friction coefficient is constant and the viscous friction coefficient is correctly estimated with $\tilde{f}^{\mathrm{v}}$, Wilk’s theorem shows that $\text{GLR}/\sigma_{v}^{2}\sim\chi^{2}$, where $\sigma_{v}^{2}$ is the variance of the noise in the signal [45]. This provides a good way to determine the threshold: if we want to have a false positive probability smaller than $10^{-9}$, we find the value of the threshold such that $P(\chi^{2}>\text{GLR}_{\text{thr}}/\sigma_{v}^{2})<10^{-9}$. However, a high threshold will also reduce the probability of detecting changepoints. An analysis of this trade-off is done in Section V.
@@
The output of the algorithm is the final list $\left\{k^{\text{cp}}_{i}\right\}_{i=1}^{N_{\text{cp}}}$ of all changepoints detected in the window of measurements.
@@ IV-B Friction Estimation
Once the changepoints have been obtained, we want to determine the dry friction coefficient between changepoints, plus the viscous friction in the data. Since the friction model we are using is linear both in the dry and the viscous friction parameters, we can use least-squares to estimate these parameters.
@@
This stage of the algorithm takes as input the original signal $X=\left\{\left(\omega_{k},f_{k}\right)\right\}_{k=1}^{N}$, together with the location of the changepoints estimated in the previous step $\left\{k_{i}^{\text{cp}}\right\}_{i=1}^{N_{\text{cp}}}$ obtained in the changepoint detection. The algorithm’s output is a vector with the value of the dry friction coefficient at each interval, and the viscous friction, which was assumed to remain constant along the whole trajectory being analyzed.
@@
First, we determine the regions in which the dry friction coefficient stays constant: the intervals between changepoints. Each interval $i$ is defined by the first ($k_{i}^{o}$) and last ($k_{i}^{f}$) measurement included in it. These measurement indexes are computed as:
$k_{i}^{\mathrm{o}}$ $=\begin{cases}0\quad&\text{if }i=1,\\ k_{i-1}^{\mathrm{cp}}+\Delta k_{\mathrm{error}}\quad&\text{otherwise},\end{cases}$ (12)
$k_{i}^{\mathrm{f}}$ $=\begin{cases}N\quad&\text{if }i=N_{\mathrm{cp}}+1,\\ k_{i}^{\mathrm{cp}}-\Delta k_{\mathrm{error}}\quad&\text{otherwise}.\end{cases}$ (13)
@@
Due to the randomness in the noise affecting the measured signals, a changepoint may be detected some steps before or after the actual change in the friction coefficient. We ignore the points near the changepoints to avoid the negative effects that these may have on estimating these coefficients. We can take $\Delta k_{\mathrm{error}}=w/2$ for small window sizes, which has been observed to give good results. With this process we obtain a total of $N_{\mathrm{int}}$ intervals, where $N_{\mathrm{int}}=N_{\mathrm{cp}}+1$.
The model in (4a) can then be written for all steps $k$ in the following linear algebraic form:
$Y=H\begin{bmatrix}F\\ f^{\mathrm{v}}\end{bmatrix}+v$ (14)
where $H$ and $Y$ are defined in (15) as the concatenation of individual $H_{i}$ and $Y_{i}$ matrices (given by (16)) for each interval, $v$ is Gaussian noise with covariance $\sigma_{v}$, and $F$ is a vector containing the dry friction coefficients for each interval:
$H=\begin{bmatrix}H_{1}\\ H_{2}\\ \vdots\\ H_{N_{\text{int}}}\end{bmatrix},\quad Y=\begin{bmatrix}Y_{1}\\ Y_{2}\\ \vdots\\ Y_{N_{\text{int}}}\end{bmatrix},\quad F=\begin{bmatrix}f_{1}^{\mathrm{d}}\\ f_{2}^{\mathrm{d}}\\ \vdots\\ f_{N_{\mathrm{int}}}^{\mathrm{d}}\end{bmatrix}$, (15)
and each individual matrix is defined as:
@@
$H_{i}$ $=\left[\begin{array}[]{ccc|c}&&&\omega_{k_{i}^{\mathrm{o}}}\\ \mathbb{0}_{\Delta k_{i}\times i}&\mathbb{1}_{\Delta k_{i}\times 1}&\mathbb{0}_{\Delta k_{i}\times(N_{\text{int}}-i)}&\vdots\\ &&&\omega_{k_{i}^{\mathrm{f}}}\\ \end{array}\right],$ (16d)
$Y_{i}$ $=\begin{bmatrix}\hat{f}_{k_{i}^{\mathrm{o}}}&\hat{f}_{k_{i}^{\mathrm{o}}+1}&\cdots&\hat{f}_{k_{i}^{\mathrm{f}}}\end{bmatrix}^{T},$ (16e)
@@
where $\Delta k_{i}=k_{i}^{\mathrm{o}}-k_{i}^{\mathrm{f}}$ is the number of points in interval $i$, and $\mathbb{0}_{a\times b}$ and $\mathbb{1}_{a\times b}$ are matrices of shape $a\times b$ filled with zeros and ones, respectively.
@@
Since Equation (14) is linear in $F$ and $f^{\mathrm{v}}$, we can use a least-squares to obtain an unbiased estimation of $F$ and $f^{\mathrm{v}}$:
$\begin{array}[]{rl}\begin{bmatrix}\hat{F}\\ \hat{f}^{\mathrm{v}}\end{bmatrix}&=\left(H^{T}H\right)^{-1}H^{T}Y.\end{array}$
@@
Due to the random nature of the noise, there is an unavoidable probability that the changepoint detector is triggered at false positives. To analyze the likelihood that a certain changepoint was incorrectly detected, we define the metric $c_{i}^{\text{rjct}}$, corresponding to the increase in the squared error of the friction estimation if the changepoint $i$ is rejected (that is, ignored), and intervals $i$ and $i+1$ are assumed to have the same friction coefficient. It is computed as:
$c_{i}^{\text{rjct}}=\frac{(H^{T}H)_{i:i}(H^{T}H)_{(i+1):(i+1)}}{(H^{T}H)_{i:i}+(H^{T}H)_{(i+1):(i+1)}}(\hat{F}_{i}-\hat{F}_{i+1})^{2},$ (17)
@@
and will be used in the next step to decide which changepoints to consider false positives.
@@
With the total dry and viscous friction torques at each interval obtained in this stage, the next stage will focus on determining the individual contribution of each FSS to the friction.

### v2

@@ III-B FSS
The FSS emulates unobservable phenomena inside the bearing of the RWA, such as the distribution of lubricant, the relative position of the inner and outer races, or any other aspect that may affect the dry friction. The state variable $q$ is the current configuration of the system and $f$ is the generated friction. The configuration represents, for instance, various stable positions of the bearings, or whether there is oil in a certain gap in the bearings or not.
Since we only use the FSS for modeling these friction phenomena, we impose some restrictions on their behavior based on observations from the TAS dataset and previous experience with friction in RWAs [26, 38]. The new configuration of the system $q$ depends only on the previous one, while its duration $\tau$ depends only on the new $q$. The transition function can then be written as
$P(q^{\prime},\tau^{\prime}|q,\tau)=P^{\mathrm{t}}(\tau^{\prime}|q^{\prime})P^{\mathrm{q}}(q^{\prime}|q)$ (6)
The transfunctions $P^{\mathrm{q}}$ and $P^{\tau}$ can be further restricted. First, we assume an FSS can only transition to adjacent configurations, that is, $P^{\mathrm{q}}(q^{\prime}|q)=0$ if $|q^{\prime}-q|\neq 1$. The index $q$ of the configuration is related to the friction: a higher $q$ implies a higher friction $f$, and there is no overlap in the range of values of friction produced by two different configurations. Figure 3 shows an example of friction probabilities, in different colors for each $q$.
To model the behavior of an RWA, we use several FSS models, each indexed by an integer $s$. Each FSS is consequently characterized by three functions: the configuration transition function $P_{s}^{\mathrm{q}}(q|q^{\prime})$ that determines how the configuration can evolve, the friction function $P_{s}^{\mathrm{f}}(f|q)$ that associates to each configuration the friction it can produce, and the duration function $P_{s}^{\mathrm{t}}(\tau|q)$ that determines how long the system stays in a given configuration.
An example of the friction produced by an RWA modeled with two FSS models is shown in Figure 4, where its similarity to the examples shown in Figure 2 can be noted.
@@
Figure 3: Evolution of the state and friction of an FSS. The left nodes show the three configurations and the transitions between them. The right side shows the friction distributions of each configuration, together with a possible trajectory of the friction value generated by the FSS as it jumps between configurations.
Example 1 (Short-term friction changes)
@@
$P_{1}^{\mathrm{f}}(f|q)$ $\sim\mathcal{U}_{\mathrm{c}}[0.3q,0.6q]$ (7a)
$P_{1}^{\mathrm{t}}(\tau|q)$ $\sim\begin{cases}\mathcal{U}_{\mathrm{d}}\{10000,20000\}\quad&\text{if }q=0\\ \mathcal{U}_{\mathrm{d}}\{0,200\}\quad&\text{if }q=1\end{cases}$ (7b)
$P_{1}(q^{\prime}|q)$ $=\mathbb{1}_{q\neq q^{\prime}}/\textstyle{\sum_{q}}\mathbb{1}_{q^{\prime}\neq q}$ (7c)
where $\mathcal{U}_{\mathrm{d}}\{a,b\}$ and $\mathcal{U}_{\mathrm{c}}[a,b]$ are uniform discrete and continuous distributions, respectively, supported in the interval $[a,b]$. The symbol $\mathbb{1}_{p}$ represents the indicator function, equal to $1$ when $p$ is true and 0 otherwise.
Example 2 (Long-term friction changes)
We want to model long-term variations of the friction switching between three configurations, with the same transition probabilities. Such variations can be due to, for example, changes in the angle of the cage inside the bearing. The probability functions are given by:
$P_{2}^{\mathrm{f}}(f|q)$ $\sim\mathcal{U}_{\mathrm{c}}[0.4q,0.6q]$ (8a)
$P_{2}^{\tau}(\tau|q)$ $\sim\mathcal{U}_{\mathrm{d}}\{10000,30000\}$ (8b)
$P_{2}(q^{\prime}|q)$ $=\mathbb{1}_{q\neq q^{\prime}}/\textstyle{\sum_{q}}\mathbb{1}_{q^{\prime}\neq q}$ (8c)
Consider the friction system proposed in Equations (4) and (5) with the FSS defined by the probability functions described above with parameters $f^{\bar{\mathrm{d}}}=f^{\mathrm{v}}=1$. The evolution of the friction components, along with the total dry friction coefficient, is shown in Figure 4.
Figure 4: Visualization of the two friction systems described in Examples 1 (FSS 1) and 2 (FSS 2). The first two plots show the evolution of the state variables of the system. The third plot shows the resulting total dry friction as in (5), and the fourth the total friction after adding a viscous component and noise, as in (4).
@@
In the paper, we assume knowledge of $q_{\text{max}}$, $P_{s}^{\tau}(\tau|q)$ and $P_{s}^{\mathrm{q}}(q|q^{\prime})$. Such information may be derived from historical data, expert knowledge, or both. We do not assume that the friction probability function $P_{s}^{\mathrm{f}}(f|q)$ is known.
@@ III-C Anomaly Modeling
To conclude this section, we define how anomalies in the RWA are represented in the proposed model. As mentioned in Section I-D, several possible anomalies are considered, and all of them affect the friction experienced in the bearing of the RWA.
We consider each anomaly to relate to one friction component. For instance, the dry friction anomaly $\theta^{\mathrm{d}}$ affects the base dry friction coefficient, while the viscous friction anomaly $\theta^{\mathrm{v}}$ affects the viscous friction coefficient.
Each FSS can be affected by one anomaly, represented by $\theta_{s}$, where $s$ is the index of the affected FSS. We assume that the effect of the anomaly manifests itself only in the friction distribution function, that is, in the values of the friction the system produces for every configuration. This relationship can be expressed as:
@@
Although the probability functions are unknown, they can be estimated based on the dataset $\mathcal{D}$.
IV Anomaly Detection Algorithm
The proposed anomaly detection algorithm uses the model developed in Section III to analyze the window of measurements under investigation and diagnoses the status of the RWA. The possible anomalies affect the friction values that each component (base dry, viscous and the FSSs) produces. The objective, then, is to extract the relevant information from the data to do the classification.
The anomaly detection algorithm comprises four stages. The first stage, changepoint detection, determines the time-steps in which an abrupt friction change has occurred. The second stage is the friction estimation, which determines the dry and viscous friction coefficients before and after every changepoint, that is, $f_{k}^{\mathrm{d}}$ and $f^{\mathrm{v}}$ in (4). The friction coefficients are used to calculate how much the friction changes at every changepoint. The third stage is the maximum likelihood assignment, which determines which FSS has produced each changepoint by maximizing the likelihood of the observed data. With this information, we can distinguish the individual contributions $f^{\bar{\mathrm{d}}}$ and $f_{k}^{s}$ in (5). Finally, once the dry friction, viscous friction, and friction produced by the switching systems is known, the anomaly classification stage produces an estimated status $\hat{\theta}$ using an SVM trained on the dataset provided. A diagram illustrating the stages of the algorithm, along with the data flow between them, is shown in Figure 5.
@@ IV Anomaly Detection Algorithm
Figure 5: Structure of the proposed anomaly detection algorithm. It consists of 4 stages. The changepoint detection determines the location of the sudden changes of friction $k_{z}^{\mathrm{jump}}$. The friction estimation estimates the friction coefficients between changepoints. The interval assignment assigns the changepoints to the different FSS in the satellite. Finally, the friction classification uses the friction values associated to the FSS to determine the anomalies present in the RWA.
@@ IV-A Changepoint Detection
First, we compute a metric that estimates the probability that the friction has changed for every step $k$ in the data. Then, we find the local maxima of this metric that are above a certain threshold. These maxima are our estimates of where the changepoints are located.
@@
with $C$ being a constant that gets canceled in the computation of the wGLR, and $W_{\mathrm{b}}$ a parameter used to penalize large deviations of the viscous friction from the values that we expect to find in the data.
@@
Similarly to the Log-Likelihood Ratio (LLR), the wGLR metric computes the log-likelihood of two hypotheses: a jump taking place at $k$, or no jump at all. One of the issues with the LLR is the need to consider the variables involved in both hypotheses, that is, the dry and viscous friction coefficients. This problem can be addressed through marginalization, leading to the Marginalized Likelihood Ratio or through maximization for the parameters, as we do in this problem, leading to the Generalized Likelihood Ratio [17].
@@
The GLR was developed for the case when there is only one changepoint to detect, but our problem contain several. Rather than computing the likelihood for the whole dataset, this issue is addressed by computing it in a window of $w$ points around the hypothetical changepoints. The constant $w$ is a hyperparameter selected from historical data, on the condition that this window size is smaller than the minimum distance between changepoints expected in the data.
@@
The last element in the metric is the use of a prior for the viscous friction $P(f^{\mathrm{v}})$. Its purpose is to avoid obtaining parameters in the optimization of (10) that fit the data well but are not realistic. As we will see in Section V-1, a high penalization constant $W_{\textrm{b}}$ can improve the algorithm’s effectiveness when the window size $w$ is small. The base viscous friction coefficient $\bar{f}^{\mathrm{v}}$ is obtained from historical data and, as it will be explained also in Section V-1, does not need to be very accurate to fulfill its objective.
@@
Once the wGLR has been computed, we must determine which points correspond to changepoints. The original GLR algorithm solves this problem by taking the point $k$ at which the metric is maximized. In our case, however, we must take into account that there are several changepoints in the complete dataset. Therefore, we first find windows of contiguous measurements where the GLR is above a threshold $\text{GLR}_{\text{thr}}$ and, for each window, we designate the maximum among those points to be a changepoint $k^{\text{cp}}$.
@@
The effectiveness of this stage of the algorithm depends on how the threshold $\text{GLR}_{\mathrm{thr}}$ is set. Under the assumption that the dry friction coefficient is constant and the viscous friction coefficient is correctly estimated with $\tilde{f}^{\mathrm{v}}$, Wilk’s theorem shows that $\text{GLR}/\sigma_{v}^{2}\sim\chi^{2}$, where $\sigma_{v}^{2}$ is the variance of the noise in the signal [47]. This provides a good way to determine the threshold: if we want to have a false positive probability smaller than $10^{-9}$, we find the value of the threshold such that $P(\chi^{2}>\text{GLR}_{\text{thr}}/\sigma_{v}^{2})<10^{-9}$. However, a high threshold will also reduce the probability of detecting changepoints, a trade-off that will be analyzed in Section VI.
@@
The output of the algorithm is the list $\left\{k^{\text{cp}}_{i}\right\}_{i=1}^{N_{\text{cp}}}$ of all changepoints detected in the window of measurements.
@@ IV-B Friction Estimation
Once the changepoints have been obtained, we determine the dry friction coefficient between changepoints, together with the viscous friction in the data. Since the friction model we are using is linear in both friction parameters, we can use least-squares to estimate these parameters.
@@
This stage of the algorithm takes the original signal $X=\left\{\left(\omega_{k},f_{k}\right)\right\}_{k=1}^{N}$, together with the changepoint locations estimated in the previous step $\left\{k_{i}^{\text{cp}}\right\}_{i=1}^{N_{\text{cp}}}$. The algorithm’s output is a vector with the value of the dry friction coefficient at each interval, together with the viscous friction, which was assumed to remain constant along the whole trajectory being analyzed.
@@
First, we determine the regions in which the dry friction coefficient stays constant, that is, the intervals between changepoints. Each interval $i$ is defined by the first ($k_{i}^{o}$) and last ($k_{i}^{f}$) measurement included in it. These measurement indexes are computed as:
$k_{i}^{\mathrm{o}}$ $=\max\left(k_{i-1}^{\mathrm{cp}},0\right)$ (12)
$k_{i}^{\mathrm{f}}$ $=\min\left(k_{i}^{\mathrm{cp}},N\right)$ (13)
@@
Due to the noise affecting the measured signals, a changepoint may be detected some steps before or after the actual change in the friction coefficient. We ignore the points near the changepoints to avoid the negative effects that these may have on estimating these coefficients. We can take $\Delta k_{\mathrm{error}}=w/2$ for small window sizes, which has been observed to give good results. With this process we obtain a total of $N_{\mathrm{int}}$ intervals, where $N_{\mathrm{int}}=N_{\mathrm{cp}}+1$.
For each interval $i$, the model in (4) can be written as
$\hat{f}_{k}$ $=f_{i}^{\mathrm{d}}+f_{v}\omega_{k}+v.$ (14)
We collect all the dry friction coefficients $f_{i}^{\mathrm{d}}$ in a vector $F$ and define the observation matrix $H_{i}$ and estimated friction vector $Y_{i}$ for each interval as follows:
$Y_{i}$ $=\begin{bmatrix}\hat{f}_{k_{i}^{o}}&\hat{f}_{k_{i}^{o}+1}&\cdots&\hat{f}_{k_{i}^{f}}\end{bmatrix}^{T},$ (15)
@@
$H_{i}$ $=\left[\begin{array}[]{ccc|c}&&&\omega_{k_{i}^{\mathrm{o}}}\\ \mathbb{0}_{\Delta k_{i}\times i}&\mathbb{1}_{\Delta k_{i}\times 1}&\mathbb{0}_{\Delta k_{i}\times(N_{\text{int}}-i)}&\vdots\\ &&&\omega_{k_{i}^{\mathrm{f}}}\\ \end{array}\right],$ (19)
$F$ $=\begin{bmatrix}f_{1}^{d}&f_{2}^{d}&\cdots&f_{N_{\mathrm{int}}}^{\mathrm{d}}\end{bmatrix}^{T},$ (20)
@@
where $\Delta k_{i}=k_{i}^{\mathrm{o}}-k_{i}^{\mathrm{f}}+1$ is the number of points in interval $i$, and $\mathbb{0}_{a\times b}$ and $\mathbb{1}_{a\times b}$ are matrices of shape $a\times b$ filled with zeros and ones, respectively. Then, (14) can be written in matrix form:
$Y_{i}$ $=H_{i}\begin{bmatrix}F^{T}&f_{v}\end{bmatrix}+v.$ (21)
We can then concatenate the observation matrices and estimated friction vectors to produce a single equation for all measurements:
$H$ $=\begin{bmatrix}H_{1}\\ H_{2}\\ \vdots\\ H_{N_{\mathrm{int}}}\end{bmatrix}\quad$ $Y$ $=\begin{bmatrix}Y_{1}\\ Y_{2}\\ \vdots\\ Y_{N_{\mathrm{int}}}\end{bmatrix}$ (22)
to give
$Y=H\begin{bmatrix}F^{T}&f^{\mathrm{v}}\end{bmatrix}^{T}+v$ (23)
@@
Since Equation (23) is linear in $F$ and $f^{\mathrm{v}}$, we use least-squares to obtain an unbiased estimate of $F$ and $f^{\mathrm{v}}$:
$\begin{array}[]{rl}\begin{bmatrix}\hat{F}^{T}&\hat{f}^{\mathrm{v}}\end{bmatrix}&=Y^{T}H\left(H^{T}H\right)^{-1}.\end{array}$
@@
There is a probability that the changepoint detector is triggered at false positives. To analyze the likelihood that a certain changepoint was incorrectly detected, we define the metric $c_{i}^{\text{rjct}}$, corresponding to the increase in the squared error of the friction estimation if the changepoint $i$ is rejected (that is, ignored), and intervals $i$ and $i+1$ are assumed to have the same friction coefficient. It is computed as:
$c_{i}^{\text{rjct}}=$ $\frac{(H^{T}H)_{i:i}(H^{T}H)_{(i+1):(i+1)}}{(H^{T}H)_{i:i}+(H^{T}H)_{(i+1):(i+1)}}(\hat{F}_{i}-\hat{F}_{i+1})^{2}$
- $\log P(\mathrm{false\ positive}),$ (24)
@@
where the probability of a false positive can be obtained from threshold $\mathrm{GLR}_{\mathrm{thr}}$ used for the changepoint detection, as explained at the end of Section IV-A. The rejection cost will be used in the next step to decide which changepoints to consider as false positives.
@@
With the total dry and viscous friction torques at each interval obtained, the next stage will focus on determining the individual contribution of each FSS to the friction.